One positive number is one-fifth of another number. The difference between the two numbers is 92, find the numbers.
Answer by ovie27(16)


There are two numbers, we will name them x and y.
1.)
y = (1/5)x
2.)
x - y = 92
3.)
x - ((1/5)x) = 92
(4/5)x = 92
x = 115
and then y = 23.
So, the two numbers are 115 and 23.